# Gaussian Processes enabled model calibration in the context of deep geological disposal

### v2

@@ 5.2 GP-based surrogate and sensitivity analysis
GP-based surrogate model: We split the dataset into training ($75\%$) and testing ($25\%$) datasets, which results in $150$ model realizations for training and $50$ for testing the accuracy of the surrogate model. The selected amount of training data results from an investigation of the relation between the amount of training data and the accuracy of the surrogate model prediction. On the one hand, we experienced that fewer model realizations lead to a significant decrease in the $R^{2}$ coefficient for the testing dataset. On the other hand, if more model realizations were used, only a marginal increase in the $R^{2}$ coefficient would be achieved compared to the computational cost required to train the surrogate model. Note that the number of model realizations can be further reduced and optimized using adaptive sampling strategies, see [67] for a recent review. The generation of the GP-based surrogate model with $7$ parameters and $150$ training data samples takes $12.3$ seconds.
@@
Figure 6: Accuracy evaluation of the GP-based surrogate model. The true versus predicted outputs for $5$ out of the $40$ GPs that constitute the surrogate model, for both vertical and horizontal convergence at different time instances. Each GP approximates the convergence at a different time instance as indicated in the title of each subplot. The blue dots represent training data, and the red dots represent testing data. The black diagonal line denotes the line of perfect prediction.
@@
Figure 7: Time-dependent global sensitivity analysis. First and total order Sobol’ indices over time for a vertical and b horizontal convergences, showing the influence of the primary creep parameters ($\eta_{p}$, $\sigma_{0,eq,p}$, $p_{p}$, $E_{p}$) and secondary creep parameters ($\eta_{s}$, $\sigma_{0,eq,s}$, $p_{s}$), presented from top to bottom, respectively.
@@
Figure 8: Aggregated global sensitivity analysis. Normalized first-order (a - b) and total-order (c - d) Sobol’ indices for vertical (a - c) and horizontal (b - d) convergences, showing the influence of the primary creep parameters ($\eta_{p}$, $\sigma_{0,eq,p}$, $p_{p}$, $E_{p}$) and secondary creep parameters ($\eta_{s}$, $\sigma_{0,eq,s}$, $p_{s}$), from left to right, respectively. The cumulated first-order and total Sobol’ indices are normalized against their respective highest overall values for fair comparison with the other indicators.
Convergence analysis of the Sobol’ indices was performed as specified in Section 3.3. The process begins with an initial set of $1282^{7}=$ 128 samples and iteratively increases the sample size by $2^{7}$ in each step until either a maximum of $655362^{16}=$ 65536 samples is reached or convergence is achieved. From (3.8), it was concluded that $163842^{14}=$ 16384 samples are enough to assure Sobol’ indices analysis convergence according to criterion (3.8). In particular, the maximum changes for S1 and ST were $0.0093$ and $0.0071$, respectively. The simulation time required for $2^{14}$ surrogate model evaluations was $4.5$ minutes.
@@ 5.3 Model calibration from in-situ monitoring data
The final task is to calibrate the four sensitive model parameters comprised in $\bm{\theta}$ using real drift convergence monitoring data. To achieve this, we embed the surrogate model trained in Section 5.2 into the optimization problem (3.2) using only the mean values of the GP-based surrogate prediction as defined in (4.5). Figure 9 displays the results obtained from the optimization problem. The results show that the simulation data generated from the high-fidelity model and the predictions from the GPs almost overlap, demonstrating the effectiveness of the surrogate model in calibration. It is crucial to highlight that model calibration is often a computationally demanding task, typically requiring multiple calls of the model. The global optimization of the remaining $4$ model parameters converged within $29$ iterations, which required $1805$ surrogate model evaluations. For comparison, training the GP-based surrogate model, used for both sensitivity analysis and calibration, only required $150$ model evaluations. The optimization took less than $1$ second to be completed on a standard laptop. This represents a dramatic reduction in the computational resources required to calibrate the deep-repository model, underscoring the efficiency of our surrogate modeling approach. The predicted parameter values agree very well with expert intuition so that the result can be considered reasonable.
6 Conclusion and Outlook
Deep geological disposal of hazardous materials requires robust numerical models to ensure long-term safety and stability. The calibration of such models with real-world monitoring data is essential for accurately reflecting in-situ conditions and enhancing repository management. This study contributes to advancing digital twinning for deep geological disposals by automating the calibration process using Gaussian Process (GP)-based surrogate models.
The presented workflow combines sensitivity analysis, surrogate modeling, and optimization to enable efficient calibration of a mechanical model representing the behavior of an emplacement drift in rock salt formations located in the northern main drift of Gorleben, Germany. Sensitivity analysis was conducted using time-dependent Sobol’ indices to identify key material parameters, followed by calibration of the sensitive parameters using 14 years of real-world monitoring data from the Gorleben site.
The results demonstrate the efficiency and accuracy of the proposed approach. First, a sensitivity analysis with $16,384$ surrogate model calls was performed within $270$ seconds to identify four relevant material parameters. Afterwards, the GP-based surrogate model was calibrated based on convergence measurements, including the convergence of the global optimization in $29$ iterations, $1805$ model evaluations for gradient construction, and a duration of less than $1$ second.
@@ Appendix B Appendix - Material parameter
$\rho$ Density $2200\cdot 10^{-6}$ $\textit{Gg}/\textit{m}^{3}$
[…]
@@ Acknowledgments
We are grateful for the provision and preparation of monitoring data from the Federal Company for Radioactive Waste Disposal (BGE mbH).
@@
All source code used for the presented benchmark studies will be published at zenodo.org upon acceptance of this manuscript. The data that support the findings of this study are available from the Federal Company for Radioactive Waste Disposal (BGE mbH). Restrictions apply to the availability of these data, which were used under licence for this study. Data are available from the authors with the permission of the Federal Company for Radioactive Waste Disposal (BGE mbH).

### v3

@@ 5.2 GP-based surrogate and sensitivity analysis
Figure 5: Accuracy evaluation of the GP-based surrogate model. The true versus predicted outputs for $5$ out of the $40$ GPs that constitute the surrogate model, for both a. vertical and b. horizontal convergence at different time instances. Each GP approximates the convergence at a different time instance as indicated in the title of each subplot. The blue dots represent training data, and the red dots represent testing data. The black diagonal line denotes the line of perfect prediction.
GP-based surrogate model: We split the dataset into training ($75\%$) and testing ($25\%$) datasets, which results in $150$ model realizations for training and $50$ for testing the accuracy of the surrogate model. The selected amount of training data results from an investigation of the relation between the amount of training data and the accuracy of the surrogate model prediction. On the one hand, we experienced that fewer model realizations lead to a significant decrease in the $R^{2}$ coefficient for the testing dataset. On the other hand, if more model realizations were used, only a marginal increase in the $R^{2}$ coefficient would be achieved compared to the computational cost required to train the surrogate model. Note that the number of model realizations can be further reduced and optimized using adaptive sampling strategies, see [69] for a recent review. The generation of the GP-based surrogate model with $7$ parameters and $150$ training data samples takes $12.3$ s.
@@
Figure 6: Time-dependent global sensitivity analysis. First and total order Sobol’ indices over time for a. vertical and b. horizontal convergences, showing the influence of the primary creep parameters ($\eta_{p}$, $\sigma_{0,eq,p}$, $p_{p}$, $E_{p}$) and secondary creep parameters ($\eta_{s}$, $\sigma_{0,eq,s}$, $p_{s}$), presented from top to bottom, respectively.
@@
Convergence analysis of the Sobol’ indices was performed as specified in Section 3.3. The process begins with an initial set of $1282^{7}=$ 128 samples and iteratively increases the sample size by $2^{7}$ in each step until either a maximum of $655362^{16}=$ 65536 samples is reached or convergence is achieved. From (3.8), it was concluded that $163842^{14}=$ 16384 samples are enough to assure Sobol’ indices analysis convergence according to criterion (3.8). In particular, the maximum changes for S1 and ST were $0.0093$ and $0.0071$, respectively. The simulation time required for $2^{14}$ surrogate model evaluations was $4.5$ min.
Figure 7: Aggregated global sensitivity analysis. Normalized first-order (a. - b.) and total-order (c. - d.) Sobol’ indices for vertical (a. - c.) and horizontal (b. - d.) convergences, showing the influence of the primary creep parameters ($\eta_{p}$, $\sigma_{0,eq,p}$, $p_{p}$, $E_{p}$) and secondary creep parameters ($\eta_{s}$, $\sigma_{0,eq,s}$, $p_{s}$), from left to right, respectively. The cumulated first-order and total Sobol’ indices are normalized against their respective highest overall values for fair comparison with the other indicators.
@@ 5.3 Model calibration from in-situ monitoring data
The final task is to calibrate the four sensitive model parameters comprised in $\bm{\theta}$ using real drift convergence monitoring data. To achieve this, we embed the surrogate model trained in Section 5.2 into the optimization problem (3.2) using only the mean values of the GP-based surrogate prediction as defined in (4.5). Figure 8 displays the results obtained from the optimization problem. The results show that the simulation data generated from the high-fidelity model and the predictions from the GPs almost overlap, demonstrating the effectiveness of the surrogate model in calibration. It is crucial to highlight that model calibration is often a computationally demanding task, typically requiring multiple calls of the model. The global optimization of the remaining $4$ model parameters converged within $29$ iterations, which required $1805$ surrogate model evaluations. For comparison, training the GP-based surrogate model, used for both sensitivity analysis and calibration, only required $150$ model evaluations. The optimization took less than $1$ s to be completed on a standard laptop. This represents a dramatic reduction in the computational resources required to calibrate the deep-repository model, underscoring the efficiency of our surrogate modeling approach. The predicted parameter values agree very well with expert intuition so that the result can be considered reasonable.
@@
6 Conclusions
Deep geological disposal of hazardous materials requires robust numerical models to ensure long-term safety and stability. The calibration of such models with real-world monitoring data is essential for accurately reflecting in-situ conditions and enhancing repository management. This study contributes to advancing digital twinning for deep geological disposals by automating the calibration process using Gaussian Process (GP)-based surrogate models. The presented workflow combines sensitivity analysis, surrogate modeling, and optimization to enable efficient calibration of a mechanical model representing the behavior of an emplacement drift in rock salt formations located in the northern main drift of the Gorleben salt dome in Germany.
The results demonstrate the efficiency and accuracy of the proposed approach. Initially, training of the GP-based surrogate model with $7$ input parameters, $150$ training data samples and $80$ outputs took $12.3$ s. The subsequent accuracy evaluation yielded $R^{2}$ scores between $0.960$ and $0.983$. Afterwards, a sensitivity analysis using time-dependent Sobol’ indices was performed with $16,384$ surrogate model calls within $4.5$ min to identify four relevant material parameters. Finally, the GP-based surrogate model was calibrated based on $14$ yr of convergence measurements, including the convergence of the global optimization in $29$ iterations, $1805$ model evaluations for gradient construction, and a duration of less than $1$ s. The surrogate model prediction provided both very good agreement with the monitoring data and valid values for parameters of the constitutive model TUBSsalt.
@@ Appendix B Appendix - Material parameter
$\rho$ Density $2200\cdot 10^{-6}$ Gg/m3
$Q$ Activation energy $22000$ J/mol
@@ Acknowledgments
We are grateful for the provision and preparation of monitoring data from the Bundesgesellschaft für Endlagerung mbH (BGE).
@@
All source code and simulation data used for the presented benchmark studies are published at https://doi.org/10.5281/zenodo.15049774. Restrictions apply to the availability of the monitoring data, which were used under license for this study. Monitoring data are available from the authors with the permission of the Bundesgesellschaft für Endlagerung mbH (BGE).